# On the stable Hopf invariant

### 09 Mar 2026 v1

@@ Abstract.
We provide a simplified approach to the the stable Hopf invariant. We provide short elementary proofs of the Cartan Formula, the Composition Formula, and the Transfer formula. In addition, when $\pi$ is a discrete group, we show how to extend these results to the stable category of $\pi$-spaces.
@@ 1. Introduction
Heinz Hopf introduced what is now called the Hopf invariant in his 1931 work on maps between spheres as a way to distinguish homotopy classes of maps that are invisible to homology [7]. In the 1950s, this notion was generalized by James [8, pp. 192–193], who defined invariants associated to desuspending maps of based spaces $\Sigma A\to\Sigma B$. For each positive integer $n$, the $n$th James–Hopf invariant is an operation on homotopy classes
@@
I am not aware of any place in the literature where the above results appear in the current form. However, with some effort they can be extracted from the properties of the geometric Hopf invariant of Crabb and Ranicki [4].
There are some key differences between our approach and that of Crabb and Ranicki: They fix a $\mathbb{Z}_{2}$-representation $V$ and define the geometric Hopf invariant as a natural transformation
@@
By constrast, the source and target of the operation $H$ do not involve a choice of representation and are expressed in terms of unequivariant stable homotopy classes. Even though $H$ is phrased in non-equivariant language, its actual construction requires passing through the $\mathbb{Z}_{2}$-equivariant stable category. The operations $H$ and $h_{V}$ are related by stabilization by representations and the application of a transfer map.
@@ 3. The tom Dieck splitting for $\mathbb{Z}_{2}$
Let $X$ be free unbased $\mathbb{Z}_{2}$-CW complex. We let $X/\mathbb{Z}_{2}$ be the orbit space. Then one has an Adams isomorphism [9, §II.7]
@@ 7. Proof of Addendum D
Let $T_{\ast}(\pi)$ be the category based left $\pi$-spaces and $\pi$-equivariant maps; this category has a zero object, i.e., the space $\ast$ consisting of a single point. Declare a map to be a weak equivalence (fibration) if and only if it is a weak equivalence (resp. fibration) upon application of the forgetful functor $T_{\ast}(\pi)\to T$. A map is a cofibration if it possesses the left lifiting property with respect to the acyclic fibrations. Then with respect to these choices, $T_{\ast}(\pi)$ is a model category by application of [6, thm. 11.6.1]. Note that the cofibrant objects of $T_{\ast}(\pi)$ are retracts of objects built up from a point by attaching free cells, where a free cell of dimension $j$ is $D^{j}\times\pi$. In particular, a free $\pi$-CW complex (in the base sense) is cofibrant. Note that every object is fibrant.

### 10 Mar 2026 v2

@@ Abstract.
We provide a simplified approach to the stable Hopf invariant. We provide short elementary proofs of the Cartan Formula, the Composition Formula, and the Transfer formula. In addition, when $\pi$ is a discrete group, we show how to extend these results to the stable category of $\pi$-spaces.
@@ 1. Introduction
Heinz Hopf introduced what is now called the Hopf invariant in his 1931 work on maps between spheres as a way to distinguish homotopy classes of maps that are invisible to homology [8]. In the 1950s, this notion was generalized by James [9, pp. 192–193], who defined invariants associated to desuspending maps of based spaces $\Sigma A\to\Sigma B$. For each positive integer $n$, the $n$th James–Hopf invariant is an operation on homotopy classes
@@
For the Segal-Snaith Hopf invariant, the Cartan formula appears in [2] (for a geometric description, see [10, thm.2.2]; see also [11] for additional properties of the Segal-Snaith Hopf invariant). For the geometric Hopf invariant of Crabb and Ranicki, a version of the Transfer and Composition Formulas appear in [5, prop. 5.33].
Our approach most closely resembles the the one taken by Crabb and Ranicki with some key differences: They fix a $\mathbb{Z}_{2}$-representation $V$ and define the geometric Hopf invariant as a natural transformation
@@
By constrast, the source and target of the operation $H$ do not involve a choice of representation and are expressed in terms of unequivariant stable homotopy classes. Even though $H$ is phrased in unequivariant language, its actual construction requires passing through the $\mathbb{Z}_{2}$-equivariant stable category. The operations $H$ and $h_{V}$ are related by stabilization by representations and the application of a transfer map.
@@ 3. The tom Dieck splitting for $\mathbb{Z}_{2}$
Let $X$ be free unbased $\mathbb{Z}_{2}$-CW complex. We let $X/\mathbb{Z}_{2}$ be the orbit space. Then one has an Adams isomorphism [12, §II.7]
@@ 7. Proof of Addendum D
Let $T_{\ast}(\pi)$ be the category based left $\pi$-spaces and $\pi$-equivariant maps; this category has a zero object, i.e., the space $\ast$ consisting of a single point. Declare a map to be a weak equivalence (fibration) if and only if it is a weak equivalence (resp. fibration) upon application of the forgetful functor $T_{\ast}(\pi)\to T$. A map is a cofibration if it possesses the left lifiting property with respect to the acyclic fibrations. Then with respect to these choices, $T_{\ast}(\pi)$ is a model category by application of [7, thm. 11.6.1]. Note that the cofibrant objects of $T_{\ast}(\pi)$ are retracts of objects built up from a point by attaching free cells, where a free cell of dimension $j$ is $D^{j}\times\pi$. In particular, a free $\pi$-CW complex (in the base sense) is cofibrant. Note that every object is fibrant.